PROBLEM 4. In a frustum of right circular cone the altitude is 3 cm and the radii of the bases are 2 cm and 6 cm. Find the lengths of the slant height and altitude of the entire cone of which this frustum is a part.
FINAL ANSWER: ______________________________ ____________________________________
1
Expert's answer
2020-12-15T07:13:18-0500

H/h=R/r

(3+h)/h=6/2

3+h=3h

3=2h

h=1.5cm

H=3+1.5

the height of the cone is 4.5cm

4.52 + 62

20.25 + 36=56.252

sguare root of 56.25=7.5cm

the slant height of the cone is 7.5cm